# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@ 2.2 Scalarization Methods
If ${\bm{\theta}}^{*}$ is unique for a given ${\bm{\alpha}}$, it is considered Pareto optimal. However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator. This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18]. Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) remains non-differentiable.
3 Finding a Single Pareto Optimal Solution
In some scenarios such as multi-task learning (MTL) [3, 78, 66], it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set. Here, all objectives are usually treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\dots,\frac{1}{m}]^{\top}$), and the goal is to produce a single model that effectively balances all objectives. The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL. However, EW may cause some tasks to have unsatisfactory performance [79, 12]. Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
$\min_{{\bm{\theta}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}}),$ (5)
where $\lambda_{i}$ is the weight for the $i$-th objective $f_{i}({\bm{\theta}})$ and the uniform preference vector ${\bm{\alpha}}$ is omitted for simplicity. These methods can be categorized as loss balancing methods and gradient balancing approaches. Some representative loss/gradient balancing methods are illustrated in Figure 3 (adapted from [19, 11]).
3.1 Loss Balancing Methods
This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss [9], homeostatic uncertainty of loss [10], loss scale [11, 12], and validation loss [13, 14, 16, 15].
Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations. Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
$\lambda_{i}^{(k)}=\frac{m\exp(\omega^{(k-1)}_{i}/\gamma)}{\sum_{j=1}^{m}\exp(\omega_{j}^{(k-1)}/\gamma)},\quad\omega^{(k-1)}_{j}=\frac{f^{(k-1)}_{j}}{f^{(k-2)}_{j}},$ (6)
@@ 3.1 Loss Balancing Methods
Kendall et al. [10] propose Uncertainty Weighting (UW) that considers the homoscedastic uncertainty of each objective and optimizes objective weights together with the model parameter as follows:
$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$ (7)
@@
Impartial Multi-Task Learning (IMTL) [11] balances losses across tasks, abbreviated as IMTL-L(oss). Specifically, IMTL-L encourages all objectives to have a similar loss scale by transforming each objective $f_{i}({\bm{\theta}})$ as $e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$. Similar to UW [10], IMTL-L simultaneously learns the transformation parameter ${\bm{s}}$ and the model parameter ${\bm{\theta}}$ as follows:
$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}\right).$ (8)
Similar to IMTL-L [11], Lin et al. [12] also aim to make all objective losses have a similar scale. They achieve it by performing a logarithm transformation on each objective (i.e., $\log f_{i}({\bm{\theta}})$). Moreover, they theoretically show that the transformation in IMTL-L is equivalent to the logarithm transformation when $s_{i}$ is the exact minimizer of $\min_{s_{i}}e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$ in each iteration. Hence, the logarithm transformation can recover the transformation in IMTL-L.
Ye et al. [13, 14] propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights ${\bm{\lambda}}$ based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
$\min_{{\bm{\lambda}}}~{}\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$ (9a)
$\ ~{}\mathrm{s.t.}\ ~{}{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$ (9b)
@@
Auto-$\lambda$ [15] modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$. Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach. Thus, it is more efficient than MOML [13, 14].
Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14]. FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach [80] and then proposes a multi-gradient aggregation method to solve it. Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$. Thus, it is significantly more efficient than MOML.
Random Weighting (RW) [17] randomly samples objective weights from a distribution (such as the standard normal distribution) and normalizes them into the simplex in each iteration. Compared to EW, RW has a higher probability of escaping local minima than EW due to the randomness from objective weights, resulting in better generalization performance.
Different from the above methods which are based on linear scalarization, Lin et al. [18] propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right),$ (10)
@@
3.2 Gradient Balancing Methods
Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ be the gradient of the $i$-th objective at the $k$-th iteration. This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration. Then, the model parameter is updated via ${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta{\bm{d}}^{(k)}$, where $\eta$ is a step size. To simplify notations, we omit the superscript in this section. Compared with loss balancing methods in Section 3.1, the gradient balancing approach usually achieves better performance.
Figure 3: Illustration of the update direction ${\bm{d}}$ of several representative methods: EW, DWA [9] from loss balancing methods, MGDA [6] from gradient weighting methods, and PCGrad [33] from gradient manipulation methods. They are illustrated in a two-objective learning problem (two specific gradients are labeled as ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$).
3.2.1 Gradient Weighting Methods
@@
Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [73], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses. This is formulated as:
$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
@@
${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$ (13)
where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun [6] use the Frank-Wolfe algorithm [81] to solve problem (13).
Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant. Specifically, ${\bm{d}}$ is computed by solving the following problem:
$\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~{}~{}\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$ (14)
@@
${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$ (15)
where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate the update direction ${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}$. Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
MGDA [6] and its variant CAGrad [19] can converge to a Pareto stationary point in the deterministic setting with full-gradient computations. However, in deep learning, where stochastic gradients are used, many works [20, 21, 23, 22, 24, 25, 26] explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
IMTL-G(rad) [11] aims to find ${\bm{d}}$ that has equal projections on all objective gradients. Thus, we have:
${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\quad 2\leq i\leq m,$ (16)
@@
${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (17)
@@
Nash-MTL [27] formulates problem (11) as a bargaining game [82, 83], where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$. Hence, ${\bm{d}}$ is computed by solving the following problem:
$\max_{{\bm{d}}}\sum_{i=1}^{m}\log\left({{\bm{g}}_{i}}^{\top}{\bm{d}}\right).$ (18)
@@
FairGrad [28] extends Nash-MTL [27] by computing ${\bm{\lambda}}$ through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant. Unlike Nash-MTL [27], FairGrad treats the problem as a constrained nonlinear least squares problem. Similarly, UPGrad [29] calculates ${\bm{\lambda}}$ by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}{\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}$, a quadratic programming (QP) problem.
Aligned-MTL [30] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$. Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed. Then, the eigenvalues $\{\sigma_{1},\dots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained. Finally, ${\bm{\lambda}}$ is computed by:
${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$ (20)
@@
Gradient Normalization (GradNorm) [31] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have similar magnitudes. It updates ${\bm{\lambda}}$ by solving the following problem via one-step gradient descent:
$\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$ (21)
@@
Dual-Balancing Multi-Task Learning (DB-MTL) [12] also normalizes all objective gradients to the same magnitude. ${\bm{\lambda}}$ is computed as $\lambda_{i}=\nicefrac{{\gamma}}{{\left\|{\bm{g}}_{i}\right\|}}$, where $\gamma=\max_{i\in[m]}\left\|{\bm{g}}_{i}\right\|$ is a scaling factor controlling the update magnitude. Thus, compared with GradNorm [31], this method guarantees all objective gradients have the same norm in each iteration. Moreover, they observe that the choice of the update magnitude $\gamma$ significantly affects performance.
3.2.2 Gradient Manipulation Methods
To address the conflicting gradient problem, gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
Projecting Conflicting Gradients (PCGrad) [33] considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$. Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients. Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$. Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as:
$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$ (22)

### v3

@@ 2.2. Scalarization Methods
If $\bm{\theta}$$^{*}$ is unique for a given $\bm{\alpha}$}, it is considered Pareto optimal. However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator. This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance (Lin et al., 2024b). Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) is still non-differentiable.
3. Finding a Single Pareto Optimal Solution
In some scenarios such as multi-task learning (MTL) (Zhang and Yang, 2022; Ruder, 2017; Crawshaw, 2020), it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set. Here, all objectives are usually treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\dots,\frac{1}{m}]^{\top}$), and the goal is to produce a single model that effectively balances all objectives. The most straightforward method is linear scalarization (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) (Zhang and Yang, 2022) in MTL. However, EW may cause some tasks to have unsatisfactory performance (Standley et al., 2020; Lin et al., 2023). Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
(5) $\min_{{\bm{\theta}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}}),$
where $\lambda_{i}$ is the weight for the $i$-th objective $f_{i}({\bm{\theta}})$ and the uniform preference vector $\bm{\alpha}$} is omitted for simplicity. These methods can be categorized as loss balancing methods and gradient balancing approaches. Some representative loss/gradient balancing methods are illustrated in Figure 4 (adapted from (Liu et al., 2021b, a)).
3.1. Loss Balancing Methods
This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss (Liu et al., 2019a), homeostatic uncertainty of loss (Kendall et al., 2018), loss scale (Liu et al., 2021a; Lin et al., 2023), and validation loss (Ye et al., 2021, 2024b, 2024a; Liu et al., 2022).
Dynamic Weight Average (DWA) (Liu et al., 2019a) estimates each objective weight as the ratio of the training losses in the last two iterations. Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
(6) $\lambda_{i}^{(k)}=\frac{m\exp(\omega^{(k-1)}_{i}/\gamma)}{\sum_{j=1}^{m}\exp(\omega_{j}^{(k-1)}/\gamma)},\quad\omega^{(k-1)}_{j}=\frac{f^{(k-1)}_{j}}{f^{(k-2)}_{j}}$,
@@ 3.1. Loss Balancing Methods
Kendall et al. (Kendall et al., 2018) propose Uncertainty Weighting (UW) that considers the homoscedastic uncertainty of each objective and optimizes objective weights together with the model parameter as follows:
(7) $\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$
@@
Impartial Multi-Task Learning (IMTL) (Liu et al., 2021a) balances losses across tasks, abbreviated as IMTL-L(oss). Specifically, IMTL-L encourages all objectives to have a similar loss scale by transforming each objective $f_{i}({\bm{\theta}})$ as $e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$. Similar to UW (Kendall et al., 2018), IMTL-L simultaneously learns the transformation parameter $\bm{s}$} and the model parameter $\bm{\theta}$} as follows:
(8) $\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}\right).$
Similar to IMTL-L (Liu et al., 2021a), Lin et al. (Lin et al., 2023) also aim to make all objective losses have a similar scale. They achieve it by performing a logarithm transformation on each objective (i.e., $\log f_{i}({\bm{\theta}})$). Moreover, they theoretically show that the transformation in IMTL-L is equivalent to the logarithm transformation when $s_{i}$ is the exact minimizer of $\min_{s_{i}}e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$ in each iteration. Hence, the logarithm transformation can recover the transformation in IMTL-L.
Ye et al. (Ye et al., 2021, 2024b) propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights $\bm{\lambda}$} based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
(9a) $\min_{{\bm{\lambda}}}~\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$
(9b) $\ ~\mathrm{s.t.}\ ~{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$
@@
Auto-$\lambda$ (Liu et al., 2022) modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$. Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach. Thus, it is more efficient than MOML (Ye et al., 2021, 2024b).
FORUM (Ye et al., 2024a) is proposed to improve the efficiency of MOML (Ye et al., 2021, 2024b). FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach (Liu et al., 2021c) and then proposes a multi-gradient aggregation method to solve it. Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$. Thus, it is significantly more efficient than MOML.
Random Weighting (RW) (Lin et al., 2022c) randomly samples objective weights from a distribution (such as the standard normal distribution) and normalizes them into the simplex in each iteration. RW has a higher probability of escaping local minima than EW due to the randomness from objective weights, resulting in better generalization performance.
Different from the above methods based on linear scalarization, Lin et al. (Lin et al., 2024b) propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
(10) $\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\frac{\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})}{\mu}\right),$
@@
3.2. Gradient Balancing Methods
Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ be the gradient of the $i$-th objective at the $k$-th iteration. This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration. Then, the model parameter is updated via ${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta{\bm{d}}^{(k)}$, where $\eta$ is a step size. To simplify notations, we omit the superscript in this section. Compared with loss balancing methods in Section 3.1, gradient balancing approaches usually achieve better performance.
Figure 4. Illustration of the update direction $\bm{d}$} of several representative methods: EW, DWA (Liu et al., 2019a) from loss balancing methods, MGDA (Sener and Koltun, 2018) from gradient weighting methods, and PCGrad (Yu et al., 2020a) from gradient manipulation methods. They are illustrated in a two-objective learning problem (two specific gradients are labeled as ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$).
3.2.1. Gradient Weighting Methods
@@
Sener and Koltun (Sener and Koltun, 2018) apply Multiple-Gradient Descent Algorithm (MGDA) (Désidéri, 2012), which aims to find a direction $\bm{d}$} at each iteration to maximize the minimal decrease across the losses. This is formulated as:
(12) $\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$
@@
(13) ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$
where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun (Sener and Koltun, 2018) use the Frank-Wolfe algorithm (Jaggi, 2013) to solve problem (13).
Conflict-Averse Gradient descent (CAGrad) (Liu et al., 2021b) improves MGDA (Sener and Koltun, 2018) by constraining the aggregated gradient $\bm{d}$} to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant. Specifically, $\bm{d}$} is computed by solving the following problem:
(14) $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~~\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$
@@
(15) ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$
where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate the update direction ${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}$. Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA (Sener and Koltun, 2018) when $c\rightarrow+\infty$.
MGDA (Sener and Koltun, 2018) and its variant CAGrad (Liu et al., 2021b) can converge to a Pareto stationary point in the deterministic setting with full-gradient computations. However, in deep learning, where stochastic gradients are used, many works (Liu and Vicente, 2021; Zhou et al., 2022; Xu et al., 2025b; Fernando et al., 2023; Chen et al., 2023a; Xiao et al., 2023; Zhang et al., 2025) explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
IMTL-G(rad) (Liu et al., 2021a) aims to find $\bm{d}$} that has equal projections on all objective gradients. Thus, we have:
(16) ${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\quad 2\leq i\leq m,$
@@
(17) ${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$
@@
Nash-MTL (Navon et al., 2022) formulates problem (11) as a bargaining game (Nash, 1953; Thomson, 1994), where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$. Hence, $\bm{d}$} is computed by solving the following problem:
(18) $\max_{{\bm{d}}}\sum_{i=1}^{m}\log\left({{\bm{g}}_{i}}^{\top}{\bm{d}}\right).$
@@
FairGrad (Ban and Ji, 2024) extends Nash-MTL (Navon et al., 2022) by computing $\bm{\lambda}$} through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant. Unlike Nash-MTL (Navon et al., 2022), FairGrad treats the problem as a constrained nonlinear least squares problem. Similarly, UPGrad (Quinton and Rey, 2024) calculates $\bm{\lambda}$} by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}({\bm{G}}^{\top}{\bm{G}}){\bm{\lambda}}$, a quadratic programming (QP) problem.
Aligned-MTL (Senushkin et al., 2023) considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of $\bm{G}$}. Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed. Then, the eigenvalues $\{\sigma_{1},\dots,\sigma_{R}\}$ and eigenvectors $\bm{V}$} of $\bm{C}$} are obtained. Finally, $\bm{\lambda}$} is computed by:
(20) ${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$
@@
Gradient Normalization (GradNorm) (Chen et al., 2018) aims to learn $\bm{\lambda}$} so that the scaled gradients have similar magnitudes. It updates $\bm{\lambda}$} by solving the following problem via one-step gradient descent:
(21) $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$
@@
Dual-Balancing Multi-Task Learning (DB-MTL) (Lin et al., 2023) also normalizes all objective gradients to the same magnitude. $\bm{\lambda}$} is computed as $\lambda_{i}=\nicefrac{{\gamma}}{{\left\|{\bm{g}}_{i}\right\|}}$, where $\gamma=\max_{i\in[m]}\left\|{\bm{g}}_{i}\right\|$ is a scaling factor controlling the update magnitude. Thus, compared with GradNorm (Chen et al., 2018), this method guarantees all objective gradients have the same norm in each iteration. Moreover, they observe that the choice of the update magnitude $\gamma$ significantly affects performance.
3.2.2. Gradient Manipulation Methods
To address gradient conflicting , gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020a) considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$. Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients. Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$. Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as:
(22) $\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$